OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (-3, -5) and (-11, -7)? is it (-4, -6)??

jimthompson5910 (jim_thompson5910):

Add the x coordinates to get ?

jimthompson5910 (jim_thompson5910):

you somehow did -11 + 3 = -8/2 = -4, which isn't correct

OpenStudy (anonymous):

=[ (-7, -6)?

OpenStudy (anonymous):

yes you got it

jimthompson5910 (jim_thompson5910):

very nice work